OpenStudy (anonymous):

Diyadiya:

OpenStudy (anonymous):

\[\Large f(x) = \left\{ {\begin{array}{*{20}{c}} {2x + 3}&{x \le 2}\\ {2x - 3}&{x > 2} \end{array}} \right.\]

OpenStudy (anonymous):

Latex: f(x) = \left\{ {\begin{array}{*{20}{c}} {2x + 3}&{x \le 2}\\ {2x - 3}&{x > 2} \end{array}} \right.

OpenStudy (anonymous):

Eh doesn't show the code properly :'(

OpenStudy (diyadiya):

ok wait

OpenStudy (diyadiya):

\[\huge \begin{cases} \frac{|x|}{x} & \text{ if } x\neq 0 \\ 0 & \text{ if } x=0 \end{cases}\]

OpenStudy (diyadiya):

Lol

OpenStudy (anonymous):

I use MathType to type up my equations, then it automatically generates the Latex code for me =)

OpenStudy (diyadiya):

\[\huge ThankYou :)\]